I'm not sure if the problem should go here but here goes...
The volume of a box with a rectangular base is 300cm cubed. The lengths of the base are 6cm and 10cm. I need to find the height of the box and the sufrace area of the box. I tried 6*10*h=300 but it didn't work. I wanted to find the area of the base in hope that it would help me but I don't know what to do after. Please help!
Thanks!

Answer by venugopalramana(3286)

I'm not sure if the problem should go here but here goes...
The volume of a box with a rectangular base is 300cm cubed. The lengths of the base are 6cm and 10cm. I need to find the height of the box and the sufrace area of the box. I tried 6*10*h=300 but it didn't work.
WHY NOT ?IT IS PERFECTLY CORRECT..CONTINUE..
60H=300
H=300/60=5 CM.
SURFACE AREA OF BOX HAS 6 FACES ALL RECTANGLES..2 WITH 6x10=120 CM^2
2 WITH 6x5 FACES=60 CM^2...AND
2 WITH 5x10 FACES=100 CM^2..
SO TOTAL SURFACE AREA =120+60+100=280 CM^2
THE FORMULA IS 2(LB+BH+HL) WHERE L,B,H ARE LENGTH ,BREADTH AND HEIGHT OF THE BOX.
I wanted to find the area of the base in hope that it would help me but I don't know what to do after. Please help!